OpenStudy (anonymous):

Will someone help me please Variable y varies directly with x2, and y = 180 when x = 6. Enter the constant of variation in the quadratic variation equation.

OpenStudy (anonymous):

@jojo4eva

OpenStudy (jhannybean):

direct variation:: \(y=kx^2\) Rearrange to solve for k: \(k = \dfrac{y}{x^2}~,~ y = 180~,~ x=6\)

OpenStudy (anonymous):

6